# Iterative Inversion of Ill-Conditioned MIMO Channels Using Symmetric Rank-$1$ Regularization

### v3

@@ Abstract
While iterative matrix inversion methods excel in computational efficiency, memory optimization, and support for parallel and distributed computing when managing large matrices, their limitations are also evident in multiple-input multiple-output (MIMO) fading channels. These methods encounter challenges related to slow convergence and diminished accuracy, especially in ill-conditioned scenarios, hindering their application in future MIMO networks such as extra-large aperture array. To address these challenges, this paper proposes a novel matrix regularization method termed symmetric rank-$1$ regularization (SR-$1$R). The proposed method functions by augmenting the channel matrix with a symmetric rank-$1$ matrix, with the primary goal of minimizing the condition number of the resultant regularized matrix. This significantly improves the matrix condition, enabling fast and accurate iterative inversion of the regularized matrix. Then, the inverse of the original channel matrix is obtained by applying the Sherman-Morrison transform on the outcome of iterative inversions. Our eigenvalue analysis unveils the best channel condition that can be achieved by an optimized SR-$1$R matrix. Moreover, a power iteration-assisted (PIA) approach is proposed to find the optimum SR-$1$R matrix without need of eigenvalue decomposition. The proposed approach exhibits logarithmic algorithm-depth in parallel computing for MIMO precoding. Finally, computer simulations demonstrate that SR-$1$R has the potential to reduce the required iteration by up to $35\%$ while achieving the performance of regularized zero-forcing.
Index Terms:
Ill-conditioned channel, iterative matrix inversion, multiple-input multiple-output (MIMO), symmetric rank-$1$ regularization (SR-$1$R).
@@ I Introduction
Iterative matrix inversion methods play a pivotal role in scenarios where direct methods become impractical or computationally expensive. This significance becomes especially pronounced when handling large matrices, as iterative techniques offer superior computational and memory efficiency, enhanced numerical stability, and support for parallel and distributed computing [2, 3, 4]. These attributes distinctly position iterative methods as advantageous over direct inversion techniques, as already demonstrated by many studies of massive multiple-input multiple-output (MIMO) systems [5, 6, 7, 8, 9].
Despite their promise, current iterative methods face difficulties with slow convergence and reduced accuracy as the wireless channel becomes ill-conditioned in future multi-user MIMO systems. Driven by the ever growing demand for spectral efficiency, user terminals are equipped with more antennas; this is pushing multi-user MIMO back to (near) symmetric architecture where the channel is extremely ill-conditioned [10]. With the increase of user antennas, increasing the service antennas to maintain the service-to-user antenna ratio as in massive MIMO systems becomes hardware costing and is challenging to device manufactory for the calibration of antenna array [11]. Moreover, when the operating frequency is fixed, adding more antennas often means enlarging the aperture of antenna-array. This could drag user terminals into the near field of the transmitter and transform conventional MIMO systems into an extra-large aperture array (ELAA) MIMO (see [12, 13, 14] for the concept). The ELAA-MIMO features significant channel spatial non-stationarity on the network side and exacerbates the ill-conditioning of wireless channel beyond what is typical in conventional MIMO systems [15, 16, 17, 18, 19, 20, 21, 22]. Notably, these challenges are not unique to iterative matrix inversion methods. Linear system inverse approaches like the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm, sketch-and-projection, sidestepping matrix inversion, also confront similar issues [7, 9, 6, 8].
@@
This paper introduces a novel channel matrix regularization approach, termed symmetric rank-$1$ regularization (SR-$1$R). As per the block diagram illustrated in Fig. 1, we consider a positive definite matrix $\mathbf{A}\in\mathbb{C}^{N\times N}$, where $N$ represents the matrix size. Rather than directly computing the inverse matrix $\mathbf{A}^{-1}$, the aim of SR-$1$R is to firstly find the inverse of a rank-$1$ regularized matrix $\mathbf{R}\in\mathbb{C}^{N\times N}$ defined as:
@@
[…]
@@
An intuitive explanation of the SR-$1$R is to reduce the largest eigenvalue below the second largest one while simultaneously increasing the modulus of the smallest eigenvalue above the second smallest one. This design is of interest because the largest and smallest eigenvalues can heavily influence the channel ill-conditioning. For instance, the largest eigenvalue dominates the ill-conditioning when the channel is highly correlated due to line-of-sight (LoS) links, while the smallest eigenvalue can make the channel condition unstable in (near) symmetric MIMO. Notably, the SR-$1$R design is primarily targeted toward ill-conditioned channels. In typical massive MIMO systems, where the channel condition is generally favorable, the largest and smallest eigenvalues are close to their neighbors. This limits the gain brought by the SR-$1$R. Therefore, such scenarios are outside the primary scope of this design.
Alongside the novel SR-$1$R concept, major contributions of our research include:
I-1 Feasibility and optimality study of SR-$1$R
@@ I-2 Practical approaches for optimization
To tackle this challenge, our research reveals that $(\xi,\mathbf{b})$ can be optimally established solely by discerning the largest and smallest eigenvalues of $\mathbf{A}$ and their corresponding eigenvectors. Subsequently, we employ a power iteration technique to compute the largest eigenvalue of $\mathbf{A}$ and a shifted power iteration method to compute the smallest eigenvalue, concurrently obtaining their respective eigenvectors. It is shown that sufficiently accurate estimates of eigenvalues and eigenvectors can be obtained through just one iterations.
@@ I-3 Simulations and Performance Evaluation
The SR-$1$R approach has undergone extensive simulations and performance evaluation across a range of MIMO fading channel types, encompassing identical and independently distributed (i.i.d.) symmetric MIMO Rayleigh channels, i.i.d symmetric MIMO Rician channels, LoS/non-LoS mixed ELAA-MIMO channels, and LoS-dominated ELAA-MIMO channels [15]. These channels exhibit diverse and contrasting channel conditions. Our simulation results consistently demonstrate that the SR-$1$R approach substantially enhances the channel conditions in all of these channel models. Thanks to this remarkable advantage, the SR-$1$R-based iterative channel inversion exhibits a significant performance enhancement in MIMO linear precoding, reducing the required iterations by more than $4\%$ when achieving RZF performance. In the case of the i.i.d. MIMO Rician channel, the reduction in the number of iterations even reaches $35\%$. Moreover, the SR-$1$R demonstrates outstanding compatibility with parallel computing, showing only minor changes in computation latency when the MIMO size increases dramatically.
The rest of this paper is organized as follows. Section II presents the system model and problem statement. Section III presents the concept of SR-$1$R and the optimality study. Section IV presents practical approaches to attain the SR-$1$R optimality. Section V presents simulation results and performance evaluation. Finally, conclusion and outlook are presented in Section VI.
@@ II System Model, Prior Art, and Problem Analysis
The SR-$1$R approach exhibits broad applicability across various scenarios necessitating iterative matrix inversion. However, our paper specifically concentrates on its implementation within wireless MIMO channel inversion, particularly tailored to MIMO linear precoding technology. This focus is motivated by the prevalent significance of MIMO linear precoding in advanced wireless communication systems, where channel inversion holds critical role for optimizing signal transmission efficiency and reducing interference, making it an ideal area for SR-$1$R exploration.
@@ II-A MIMO Linear Precoding
Consider MIMO downlink communications, where an access point (i.e., the transmitter) equipped with $M$ transmit-antennas sends $N$ independent data streams to $N$ $(\leq M)$ user antennas, respectively. Denote $\mathbf{s}\in\mathbb{C}^{N\times 1}$ as an information-bearing symbol block in the spatial domain. To maintain the orthogonality between data streams, an interference-rejection linear precoder, denoted as $\mathbf{W}\in\mathbb{C}^{M\times N}$, is employed at the transmitter, which yields the transmit symbol block:
@@
With the ZF precoder, data streams do not interfere with each other and can be decoded separately. On the other hand, the computation of matrix inverse $\mathbf{A}^{-1}$ costs cubic complexity and does not support well the parallel computing technology [29, 30]. This presents a formidable challenge in the design of scalable large-MIMO systems.
@@ II-B Prior Art Analysis
Our analysis of prior art identifies two main categories of iterative solutions to the inversion problem: a) iterative linear system inversion, and b) iterative matrix inversion. This subsection briefly explores both categories, laying the groundwork for advocating the necessity of employing iterative matrix inversion.
II-B1 Iterative linear system inversion
In numerous use cases, the matrix inverse problem can be circumvented by transforming it into a linear system inverse problem. Specifically for the linear ZF precoding problem (5), rather than directly computing $\mathbf{A}^{-1}$, we can find a vector $\overline{\mathbf{x}}\in\mathbb{C}^{N\times 1}$ that is related to $\mathbf{x}$ through the equation: $\mathbf{x}=\mathbf{H}^{H}\overline{\mathbf{x}}$. In line with (5), the goal is to iteratively solve the following equation (e.g., [6, 32, 33]):
@@ II-B1 Iterative linear system inversion
An analogous approach can also be found in the MIMO uplink detection (e.g., [9, 34, 18]). In the context of iterative linear system inversion, two significant issues come to the forefront: 1) the performance of iterative linear system inversion is highly susceptible to the condition of the matrix $\mathbf{A}$. As shown in Section V, when dealing with ELAA-MIMO channels, which are often ill-conditioned, these iterative algorithms are too sub-optimal; 2) some inverse problems, like MIMO non-linear precoding (see [31, 10, 35, 36, 30]) cannot be transformed into the linear system inverse problems. In these cases, matrix inversion becomes a necessary step.
@@ II-B2 Iterative matrix inversion
Various techniques are available for the iterative matrix inversion. Common methods include the Neumann series method [5] and the Schulz method [2]. The Schulz method is also known as the Hotelling-Bodewig algorithm [3] or the Newton iteration [9, 37] in the literature. Additionally, the sketch-and-projection approach has been recently adopted for matrix inversion [4]. Our state-of-the-art analysis concludes that the Neumann series method typically requires more iterations compared to the Schulz method [37]. Furthermore, the performance of the sketch-and-projection method is influenced by the randomness of the sketching process, making it less stable than the Schulz method. Therefore, in this paper, we opt for the Schulz method to evaluate the performance of iterative matrix inversion concerning the condition numbers $\kappa(\mathbf{A})$ or $\kappa(\mathbf{R})$.
@@
$\mathbf{X}_{i}$ $=\omega\mathbf{A}^{H},~{}i=0$
$\mathbf{X}_{i}$ $=2\mathbf{X}_{i-1}-\mathbf{X}_{i-1}\mathbf{A}\mathbf{X}_{i-1},~{}i>0,$ (9)
@@ II-C Research Problem and Analysis
For the iterative matrix inversion, the number of iterations should be minimized while ensuring satisfactory performance, as this is crucial for maximizing computational efficiency and accuracy. The number of iterations is strongly dependent on the condition number $\kappa(\mathbf{A})$. For instance, in the case of a unitary channel matrix with $\kappa(\mathbf{A})=1$, setting $\omega=1$ in (II-B2) immediately yields $\|\mathbf{I}-\mathbf{A}\mathbf{X}_{0}\|_{\mathrm{F}}=0$.
@@
For the pre-conditioning method, the channel matrix $\mathbf{A}$ is multiplied with a pre-conditioning matrix $\mathbf{P}^{-1}$, which yields
$\mathbf{R}_{\textsc{pre}}=\mathbf{P}^{-1}\mathbf{A}.$ (12)
Iterative inversion is performed on $\mathbf{R}_{\textsc{pre}}$, which aims to generate
$\mathbf{R}_{\textsc{pre}}^{-1}=\mathbf{A}^{-1}\mathbf{P}.$ (13)
After the iterative inversion, $\mathbf{A}^{-1}$ is computed through $\mathbf{A}^{-1}=\mathbf{R}_{\textsc{pre}}^{-1}\mathbf{P}^{-1}$.
@@
$\rm{GS}:\quad$ $\mathbf{P}=\mathbf{D}+\mathbf{};$ (15)
@@
Despite extensive prior research documented in the literature, pre-conditioning approaches have only yielded marginal performance enhancements when dealing with ill-conditioned ELAA-MIMO channels (as detailed in Section V). To further enhance both the performance and convergence in iterative matrix inversion, we introduce the SR-$1$R method.
III SR-$1$R: Feasibility and Optimality Study
The concept of SR-$1$R has already been introduced in Section I, as outlined in Eqn. (1)-(2). In this section, our objective is to address the question of whether there exists $(\xi,\mathbf{b})$ that can lead to:
  • a)
    $\kappa(\mathbf{R})<\kappa(\mathbf{A})$;
  • b)
    minimization of $\kappa(\mathbf{R})$.
This question motivates the feasibility and optimality study, relying on the assumption of availability of EVD (or SVD as appropriate). This assumption is reasonable here as it is specifically for feasibility and optimality analysis.
@@ III-A Feasibility Study
In the mathematical literature, (20) is well known as the rank-$1$ modification of symmetric eigenproblem [38, 39]. One of the significant findings from previous research can be summarized as:
@@
However, in this case, it cannot be guaranteed that the inequality $\kappa(\mathbf{\Theta})<\kappa(\mathbf{A})$ holds. Consequently, having $\xi<0$ is not conducive for the SR-$1$R approach.
@@
where ||$\cdot$|| ⋅ | stands for the absolute value of a number. Then, the condition number of $\mathbf{\Theta}$ can be expressed as
@@
$-\theta_{0}<\theta_{N-1}<-\theta_{N-2}.$ (27)
@@
We prove a sufficient condition of Theorem 1: $\theta_{N-1}$ monotonically decreases from 00 to $\infty$- ∞ as $\xi$ increases from $\xi^{\bot}$ to $\infty$, where $\xi^{\bot}$ is a positive scalar explained later in (31).
@@
The right hand side of (38) tends to $\infty$- ∞ as $\xi\to\infty$. Recall (1) that $\theta_{n,}$ n=0,⋯,N-2 is bounded. Hence, we must have $\theta_{N-1}\to-\infty$ as $\xi\to\infty$ in the left hand side. (37) is proved.
Therefore, $\theta_{N-1}$ monotonically decreases from 00 to $\infty$- ∞ as $\xi$ increases from $\xi^{\bot}$ to $\infty$, yielding Theorem 1. ∎
@@
This intermediate result ensures the feasibility of employing SR-$1$R to improve the matrix condition. Nevertheless, an in-depth study is needed to establish the optimality of SR-$1$R.

### v4

# Symmetric Rank-$1$ Regularization for Iterative Inversion of Ill-Conditioned MIMO Channels

@@ Abstract
While iterative matrix inversion methods excel in compatibility of parallel and distributed computing when managing large matrices, their limitations are also evident in multiple-input multiple-output (MIMO) fading channels. These methods encounter challenges related to slow convergence and diminished accuracy, especially in ill-conditioned scenarios, hindering their application in future MIMO networks such as extra-large aperture array. To address these challenges, this paper proposes a novel matrix regularization method termed symmetric rank-$1$ regularization (SR-11R). The proposed method functions by augmenting the channel matrix with a symmetric rank-$1$ matrix, with the primary goal of minimizing the condition number of the resultant regularized matrix. This significantly improves the matrix condition, enabling fast and accurate iterative inversion of the regularized matrix. Then, the inverse of the original channel matrix is obtained by applying the Sherman-Morrison transform on the outcome of iterative inversions. Our eigenvalue analysis unveils the best channel condition that can be achieved by an optimized SR-11R matrix. Moreover, a power iteration-assisted (PIA) approach is proposed to find the optimum SR-11R matrix without need of eigenvalue decomposition. The proposed approach exhibits logarithmic algorithm-depth in parallel computing for MIMO precoding. Finally, computer simulations demonstrate that SR-11R has the potential to reduce the required iteration by up to $35\%$ while achieving the performance of regularized zero-forcing.
@@ I Introduction
Iterative matrix inversion methods play a pivotal role in scenarios where direct methods become impractical or incurs prohibitive computational latency. This significance becomes especially pronounced when handling large matrices, as iterative techniques offer superior compatibility to parallel and distributed computing as well as enhanced numerical stability [2, 3, 4]. In addition, as future multiple-input multiple-output (MIMO) hardware increasingly incorporates high parallel computing capabilities (e.g., quantum computing [5, 6]), such compatibility will be essential for enabling scalable and efficient signal processing algorithms.
Despite their promise, current iterative matrix inversion methods face difficulties with slow convergence and reduced accuracy as the wireless channel becomes ill-conditioned in future multi-user MIMO systems. Driven by the ever growing demand for spectral efficiency, user terminals are equipped with more antennas; this is pushing multi-user MIMO back to (near) symmetric architecture where the channel is extremely ill-conditioned [7]. With the increase of user antennas, increasing the service antennas to maintain the service-to-user antenna ratio as in massive MIMO systems becomes hardware costing and is challenging to device manufactory for the calibration of antenna array [8]. Moreover, when the operating frequency is fixed, adding more antennas often means enlarging the aperture of antenna-array. This could drag user terminals into the near field of the transmitter and transform conventional MIMO systems into an extra-large aperture array (ELAA) MIMO (see [9, 10] for the concept). The ELAA-MIMO features significant channel spatial non-stationarity on the network side and exacerbates the ill-conditioning of wireless channel beyond what is typical in conventional MIMO systems [11, 12, 13, 14].
@@
This paper introduces a novel channel matrix regularization approach, termed symmetric rank-$1$ regularization (SR-11R). As per the block diagram illustrated in Fig. 1, we consider a positive definite matrix $\mathbf{A}\in\mathbb{C}^{N\times N}$, where $N$ represents the matrix size. Rather than directly computing the inverse matrix $\mathbf{A}^{-1}$, the aim of SR-11R is to firstly find the inverse of a rank-$1$ regularized matrix $\mathbf{R}\in\mathbb{C}^{N\times N}$ defined as:
@@
where $\xi\in\mathbb{R}$ is a scaling factor, $\mathbf{b}\in\mathbb{C}^{N\times 1}$ is a unitary vector (i.e., $\|\mathbf{b}\|=1$), $\cdot]^{H}$ and $\|\cdot\|$ denotes the Hermitian transpose and the Euclidean norm, respectively. Next, Sherman-Morrison transform (see [19]) is applied on $\mathbf{R}^{-1}$, resulting in:
@@
An intuitive explanation of the SR-11R is to reduce the largest eigenvalue below the second largest one while simultaneously increasing the modulus of the smallest eigenvalue above the second smallest one. This design is of interest because the largest and smallest eigenvalues can heavily influence the channel ill-conditioning. For instance, the largest eigenvalue dominates the ill-conditioning when the channel is highly correlated due to line-of-sight (LoS) links, while the smallest eigenvalue can make the channel condition unstable in (near) symmetric MIMO. Notably, the SR-11R design is primarily targeted toward ill-conditioned channels. In typical massive MIMO systems, where the channel condition is generally favorable, the largest and smallest eigenvalues are close to their neighbors. This limits the gain brought by the SR-11R. Therefore, such scenarios are outside the primary scope of this design.
Alongside the novel SR-11R concept, major contributions of our research include:
I-1 Feasibility and optimality study of SR-11R
@@ I-2 Practical approaches for optimization
To tackle this challenge, our research reveals that $(\xi,\mathbf{b})$ can be optimally established solely by discerning the largest and smallest eigenvalues of $\mathbf{A}$ and their corresponding eigenvectors. Subsequently, we employ a power iteration technique to compute the largest eigenvalue of $\mathbf{A}$ and a shifted power iteration method to compute the smallest eigenvalue, concurrently obtaining their respective eigenvectors. It is shown that sufficiently accurate estimates of eigenvalues and eigenvectors can be obtained through just one iteration.
@@ I-3 Simulations and Performance Evaluation
The SR-11R approach has undergone extensive simulations and performance evaluation across a range of MIMO fading channel types, encompassing identical and independently distributed (i.i.d.) symmetric MIMO Rayleigh channels, i.i.d symmetric MIMO Rician channels, LoS/non-LoS mixed ELAA-MIMO channels, and LoS-dominated ELAA-MIMO channels [11]. These channels exhibit diverse and contrasting channel conditions. Our simulation results consistently demonstrate that the SR-11R approach substantially enhances the channel conditions in all of these channel models. Thanks to this remarkable advantage, the SR-11R-based iterative channel inversion exhibits a significant performance enhancement in MIMO linear precoding, reducing the required iterations by more than $4\%$ when achieving RZF performance. In the case of the i.i.d. MIMO Rician channel, the reduction in the number of iterations even reaches $35\%$. Moreover, the SR-11R demonstrates outstanding compatibility with parallel computing, showing only minor changes in computation latency when the MIMO size increases dramatically.
@@ II System Model, Prior Art, and Problem Analysis
The SR-11R approach exhibits broad applicability across various scenarios necessitating iterative matrix inversion. However, our paper specifically concentrates on its implementation within wireless MIMO channel inversion, particularly tailored to MIMO linear precoding technology. This focus is motivated by the prevalent significance of MIMO linear precoding in advanced wireless communication systems, where channel inversion holds critical role for optimizing signal transmission efficiency and reducing interference, making it an ideal area for SR-11R exploration.
@@ II-A MIMO Linear Precoding
Consider MIMO downlink communications, where an access point (i.e., the transmitter) equipped with $M$ transmit-antennas sends $N$ independent data streams to NN $(\leq M)$ user antennas, respectively. Denote $\mathbf{s}\in\mathbb{C}^{N\times 1}$ as an information-bearing symbol block in the spatial domain. To maintain the orthogonality between data streams, an interference-rejection linear precoder, denoted as $\mathbf{W}\in\mathbb{C}^{M\times N}$, is employed at the transmitter, which yields the transmit symbol block:
@@
With the ZF precoder, data streams do not interfere with each other and can be decoded separately. On the other hand, the computation of matrix inverse $\mathbf{A}^{-1}$ costs cubic complexity and does not support well the parallel computing technology [21]. This presents a formidable challenge in the design of scalable large-MIMO systems.
@@ II-B Prior Art Analysis
Our analysis of prior art identifies two main categories of iterative solutions to the inversion problem: a) iterative LSI, and b) iterative matrix inversion. This subsection briefly explores both categories, laying the groundwork for advocating the necessity of employing iterative matrix inversion.
II-B1 Iterative LSI
In numerous use cases, the matrix inverse problem can be circumvented by transforming it into a LSI problem. Specifically for the linear ZF precoding problem (5), rather than directly computing $\mathbf{A}^{-1}$, we can find a vector $\overline{\mathbf{x}}\in\mathbb{C}^{N\times 1}$ that is related to $\mathbf{x}$ through the equation: $\mathbf{x}=\mathbf{H}^{H}\overline{\mathbf{x}}$. In line with (5), the goal is to iteratively solve the following equation (e.g., [23, 24, 25]):
@@ II-B1 Iterative LSI
An analogous approach can also be found in the MIMO uplink detection (e.g., [26]).
Iterative LSI approaches reduce the computational complexity of matrix–matrix multiplication to matrix-vector multiplications. Moreover, good initialization can substantially enhance convergence when the number of iterations is limited (e.g., around $10$). These properties make iterative LSI approaches particularly attractive under well-conditioned channel scenarios, such as in massive MIMO systems, as demonstrated in recent studies [26, 27].
When the channel is ill-conditioned, however, the required number of iterations increases, and the convergence speed is more critical. Iterative LSI methods generally exhibit linear convergence speed, whereas iterative matrix inversion algorithms can achieve exponential convergence []. Therefore, in our following discussions, we will focus our study on iterative matrix inversion.
@@ II-B2 Iterative matrix inversion
Various techniques are available for the iterative matrix inversion. Common methods include the Neumann series method [28] and the Schulz method [2]. The Schulz method is also known as the Hotelling-Bodewig algorithm [3] or the Newton iteration [26] in the literature. Additionally, the sketch-and-projection approach has been recently adopted for matrix inversion [4]. Our state-of-the-art analysis concludes that the Neumann series method typically requires more iterations compared to the Schulz method [29]. Furthermore, the performance of the sketch-and-projection method is influenced by the randomness of the sketching process, making it less stable than the Schulz method. Therefore, in this paper, we opt for the Schulz method to evaluate the performance of iterative matrix inversion concerning the condition numbers $\kappa(\mathbf{A})$ or $\kappa(\mathbf{R})$.
@@
$\mathbf{X}_{0}=\omega\mathbf{A}^{H};~\mathbf{X}_{i}=2\mathbf{X}_{i-1}-\mathbf{X}_{i-1}\mathbf{A}\mathbf{X}_{i-1},~i>0,$ (9)
@@ II-C Research Problem and Analysis
For the iterative matrix inversion, the number of iterations should be minimized while ensuring satisfactory performance, as this is crucial for maximizing computational efficiency and accuracy. The number of iterations is strongly dependent on the condition number $\kappa(\mathbf{A})$. For instance, in the case of a unitary channel matrix with $\kappa(\mathbf{A})=1$, setting $\omega=1$ in (9) immediately yields $\|\mathbf{I}-\mathbf{A}\mathbf{X}_{0}\|_{\mathrm{F}}=0$.
@@
For the pre-conditioning method, the channel matrix $\mathbf{A}$ is multiplied with a pre-conditioning matrix $\mathbf{P}^{-1}$, which yields $\mathbf{R}_{\textsc{pre}}=\mathbf{P}^{-1}\mathbf{A}$. Iterative inversion is performed on $\mathbf{R}_{\textsc{pre}}$, which aims to generate $\mathbf{R}_{\textsc{pre}}^{-1}=\mathbf{A}^{-1}\mathbf{P}$. After the iterative inversion, $\mathbf{A}^{-1}$ is computed through $\mathbf{A}^{-1}=\mathbf{R}_{\textsc{pre}}^{-1}\mathbf{P}^{-1}$.
@@
$\rm{GS}:\quad$ $\mathbf{P}=\mathbf{D}+\mathbf{L};$ (13)
@@
Despite extensive prior research documented in the literature, pre-conditioning approaches have only yielded marginal performance enhancements when dealing with ill-conditioned ELAA-MIMO channels (as detailed in Section V). To further enhance both the performance and convergence in iterative matrix inversion, we introduce the SR-11R method.
III SR-11R: Feasibility and Optimality Study
The concept of SR-11R has already been introduced in Section I, as outlined in Eqn. (1)-(2). In this section, our objective is to address the question of whether there exists $(\xi,\mathbf{b})$ that can lead to: a) $\kappa(\mathbf{R})<\kappa(\mathbf{A})$; b) minimization of $\kappa(\mathbf{R})$. This question motivates the feasibility and optimality study, relying on the assumption of availability of EVD (or SVD as appropriate). This assumption is reasonable here as it is specifically for feasibility and optimality analysis.
@@ III-A Feasibility Study
In the mathematical literature, (18) is well known as the rank-$1$ modification of symmetric eigenproblem [30]. One of the significant findings from previous research can be summarized as:
@@
However, in this case, it cannot be guaranteed that the inequality $\kappa(\mathbf{\Theta})<\kappa(\mathbf{A})$ holds. Consequently, having $\xi<0$ is not conducive for the SR-11R approach.
@@
where $|\cdot|$ stands for the absolute value of a number. Then, the condition number of $\mathbf{\Theta}$ can be expressed as
@@
$\theta_{0}<\theta_{N-1}<-\theta_{N-2}.$ (25)
@@
We prove a sufficient condition of Theorem 1: $\theta_{N-1}$ monotonically decreases from 0 to -$\infty$ as $\xi$ increases from $\xi^{\bot}$ to $\infty$, where $\xi^{\bot}$ is a positive scalar explained later in (29).
@@
The right hand side of (36) tends to -$\infty$ as $\xi\to\infty$. Recall (1) that $\theta_{n,}$ n=0,⋯,N-2 is bounded. Hence, we must have $\theta_{N-1}\to$-$\infty$ as $\xi\to\infty$ in the left hand side. (35) is proved.
Therefore, $\theta_{N-1}$ monotonically decreases from 0 to -$\infty$ as $\xi$ increases from $\xi^{\bot}$ to $\infty$, yielding Theorem 1. ∎
@@
This intermediate result ensures the feasibility of employing SR-11R to improve the matrix condition. Nevertheless, an in-depth study is needed to establish the optimality of SR-11R.